you have $3.70 in dimes and quarters, you have 5 more quarters than dimes, how many how each coin do you have using substitution method.
Answer by mananth(16949)

you have $3.70 in dimes and quarters, you have 5 more quarters than dimes, how many how each coin do you have using substitution method.
Let the dimes be x
The quarters will be x+5
The sum of their values = 370 cents
10x+25(x+5)=370
10x+25x+125=370
35x=370-125
x=245/35
x=7 which is the number of dimes
x+7= 7+5= 12 which is the number of quarters